# Emergent Mixture-of-Experts: Can Dense Pre-trained Transformers Benefit from Emergent Modular Structures?

### v1

@@ 4.1 Full Fine-tuning Performance
We evaluate EMoE’s ID performance on 5 GLUE tasks. According to Table 2: (1) On average, EMoE and EMoE-learn outperform other baselines. (2) Among the two methods that do not introduce additional parameters, EMoE significantly outperforms noisy tuning. (3) EMoE provides stable improvements over baselines across different settings, demonstrating its generality.
4.2 LoRA-Tuning performance
Table 3: ID and OOD performance of EMoE and baseline models. All the reported results are obtained from 3 independent experiments. OOD Metrics (averaged over 14 OOD tasks, lower is better) provide additional information for out-of-distribution generalization. The best result is highlighted in bold.
Algorithm MRPC CoLA RTE STSB SST2 QNLI QQP MNLI ID-Avg
OOD
BERT-Large (340M Parameters)
LoRA 89.97 63.40 72.92 90.51 93.16 92.20 87.21 85.40 84.35 4.86
Block 89.34 62.10 71.96 90.39 93.35 92.04 88.45 86.20 84.23(-0.12) 4.95
EMoE-block
89.77 63.25 71.60 90.31 93.69 92.09 88.08 86.21 84.38(+0.03) 5.89
EMoE-learn
89.87 64.00 71.36 90.48 93.65 92.40 87.55 85.62 84.37(+0.02) 4.66
GPT2-XL (1.5B Parameters)
LoRA 86.83 60.88 78.70 89.07 95.18 91.84 87.41 86.93 84.61 5.61
Block 86.59 61.18 79.78 89.08 95.45 91.88 87.71 86.95 84.83(+0.22) 5.13
EMoE-block
87.86 62.88 80.05 89.18 95.49 92.10 89.69 86.87 85.52(+0.91) 5.71
EMoE-learn
87.93 61.50 79.90 89.48 95.18 92.33 89.71 87.00 85.38(+0.77) 4.40
With the increasing scale of pre-trained models, parameter-efficient tuning (Houlsby et al., 2019) gets popular. Observing weaker results with LoRA-tuned ViT-Large on Domainbed compared to ViT-Base, we omit LoRA tuning on Domainbed. GLUE-X (Yang et al., 2023) is a recently introduced OOD dataset for the GLUE benchmark. Therefore, we assess EMoE’s ID performance on GLUE and OOD performance on GLUE-X with the standard LoRA tuning. Notably, LoRA weights are only added in attention, with all other weights frozen. Therefore, the only difference between EMoE and its dense counterpart is FFNs activation. For the OOD metric, we follow GLUE-X (Yang et al., 2023) and employ the Friedman rank (Friedman, 1940) $\text{rank}_{f}=\frac{1}{n}\sum_{i=1}^{n}\text{rank}_{i}$. For each method under the same backbone, $\text{rank}_{i}$ is produced based on the best result and the average result for each dataset. For the 13 OOD tasks used in GLUE-X, each method generates 26 $\text{rank}_{i}$ values. The OOD results presented in Table 3 represent the mean of all these $\text{rank}_{i}$ values. The original results of each task can be found in the Appendix A.4.1. For algorithm configurations, since GMoE copies the FFNs of the pre-trained model to form the MoEs, it is meaningless if the MoEs parameters aren’t tuned. Thus, in Table 3, we conduct experiments with LoRA tuning for GMoE and also tune the transformer block where the original FFNs is replaced. For comparison, we also (1) LoRA-tune dense model and fine-tune the transformer block at the corresponding layer (denoted “Block”); (2) LoRA-tune EMoE with the transformed block also updated (denoted “EMoE-block”).
According to Table 3, EMoE continues to enhance the downstream performance in LoRA tuning: (1) EMoE demonstrates significant enhancements compared to LoRA and doesn’t introduce additional trainable parameters or procedures. Notably, EMoE achieves comparable results with GMoE on BERT-large and outperforms it on GPT2-XL. (2) When the blocks are tuned, EMoE also brings improvements (EMoE-block vs. Block) (3) Consistent with the full-finetuning findings, EMoE exhibits higher stability than EMoE-learn and delivers superior overall results. (4) Besides vision, MoEs also improve OOD performance in language tasks (GMoE vs. Block, EMoE vs. LoRA).
5 Analysis
5.1 How does EMoE improve fine-tuning performance?
Refer to caption     Refer to caption
Figure 3: Left: ID and OOD results compared with LoRA for validating EMoE’s training & inference effects. Right: sparse activated training results with different expert selections.
EMoE achieves notable improvements in both ID and OOD scenarios. However, it is non-trivial that simply transforming the pre-trained FFNs into MoEs before fine-tuning can yield such benefits (especially in LoRA tuning where the MoEs part is totally not updated). Therefore, we investigate the reasons behind the enhanced fine-tuning performance. To decouple the impact of updating expert parameters, we focus on LoRA tuning. We use BERT-Large as the backbone model.
EMoE benefits LoRA weights learning instead of influencing inference. EMoE and dense model only differ in FFNs activations; such difference might (1) directly impact outputs during testing and (2) influence the parameter updating during training. In light of this, we propose two variants and compare them with vanilla LoRA tuning and LoRA tuning with EMoE: (a) LoRA2EMoE: splitting a LoRA-tuned dense model into its MoEs counterpart. If the test results differ from the original ones, we can infer that the sparse activation has an impact during testing. (b) EMoE2LoRA: merging the LoRA weights tuned with EMoE back into the dense. If no changes occur, it implies that the primary factor is the parameter updating. According to Figure 3 Left, doing sparse activation during testing has little impact (LoRA2EMoE). But when merging LoRA weights tuned with EMoE back into the dense, the performance remains significantly better than the vanilla LoRA tuning (EMoE2LoRA vs. LoRA) and is almost identical to EMoE. For full results, please refer to Appendix A.4 Table 11.
EMoE masks neurons with negative transfer impacts. The only difference between EMoE and vanilla LoRA tuning is EMoE blocks some activated neurons during training by $\operatorname{Top-k}$ expert selection. Based on these, we hypothesize that EMoE’s effects stem from preventing negative knowledge transfer from blocked neurons. Therefore, we investigate whether there are such negative transfers. Specifically, we study the following expert selection variants: (1) Bottom-k: select $k=16$ experts who get the lowest scores; (2) Not-top-k: select experts who are not among the top-k experts. These variants are evaluated across 6 tasks from GLUE. The averaged outcomes are in Figure 3 Right. Full results can be found in Appendix A.4 Table 12. LoRA tuning results with “Bottom-k” and “Not-top-k” expert selections are worse than vanilla LoRA tuning. One may attribute this drop to the reduced number of employed neurons. To further assess this, we compare “Top-k” with “Bottom-k”, where an equal number of neurons are used. Notably, “Bottom-k” significantly lags behind “Top-k”. This further corroborates that masked neurons have negative transfer effects.
5.2 Ablation studies
Figure 4: Left: Activations of neurons in FFNs a pre-trained ViT (Dosovitskiy et al., 2021). Right: By clustering the keys in the FFNs and rearranging the action values accordingly, activation patterns of neurons emerge. These co-activated patterns correspond to the modular structure within the pre-trained dense model. With appropriate partitioning, we can transform dense networks into modular networks.
Table 4: Clustering-based and random expert constructing results
Top-k Bottom-K
Cluster +0.92 -0.48
Random -0.11 -0.34
Expert Constructing Methods Figure 4 demonstrates that clustering can decompose modular components within the dense model. To provide further evidence that the EMoE’s improvements stem from leveraging modular features rather than just sparse activation, we compare the results of (1) clustering-based expert construction and (2) random construction within the same setting of Section 3. The relative changes in averaged outcomes compared to the dense baseline are shown in Table 4, while full results can be found in the appendix A.4, Table 12. It’s noteworthy that while cluster top-k exhibits a significant improvement over dense, random top-k is conversely worse than dense baseline. This suggests that random construction can negatively impact gating. Moreover, considering selecting weights with negative transfer under bottom-k selections, it’s observed that cluster bottom-k also achieves lower results. In summary, clustering-based methods can externalize implicit modularity within the dense model. Within suitable frameworks like MoEs, such modularity can facilitate downstream tuning.
Expert Constructing Configurations Beyond the settings detailed in the main results section, which are based on $N=64$ experts and top-k $\in\{16,32,48\}$, we also present specific scenarios where $N\in\{16,32\}$, and top-k varies within $\{2,4,8,16\}$ in Figure 5. Notably, in each of these settings, (1) EMoE consistently surpasses the dense model, illustrating its insensitivity to hyper-parameters. (2) On average, $\operatorname{avg-k}$ gating exhibits superior performance than learned gating. Though learned gating (EMoE-learn) outperforms $\operatorname{avg-k}$ gating in a few specific settings (Figure 5 (b) and (e)). This is consistent with the earlier results in Section 4. Regarding how many EMoE layers should be introduced, our findings align with those discovered in GMoE, indicating that only a limited number of layers can be converted into the EMoE layer. If excessive EMoE layers are introduced, performance deteriorates. Taking GPT2-XL (48 layers) as an example, when introducing EMoE every 2 layers in the latter half, the performance averaged across 5 GLUE tasks (79.36) matches that of the dense model (78.87). But when adopting EMoR every 2-layer for the entire model, the performance lags slightly behind that of the dense model (78.17) but surpasses EMoE-learn (75.87). For additional configurations, please refer to the appendix A.4.2 Table 14.
Figure 6: Expert selections during training with distinct gating functions ($\operatorname{avg-k}$ vs. learned) and expert types (modules from dense vs. repetitions of dense). The vertical axis illustrates training steps (top-down arrangement signifies begin-end); the horizontal axis represents expert selection frequency within 1K steps (deeper color implies a higher frequency). (a), (b) and (c) correspond to EMoE, EMoE-learn, and GMoE.
Expert Selection Changes During Training To further understand $\operatorname{avg-k}$ gating and learned gating, we visualize expert selections of GPT2-XL during fine-tuning on 6 tasks with 16 experts. In Figure 6, we showcase the results for the largest dataset QNLI among them. Full results are available in the appendix A.5. Our observations are: (1) Both $\operatorname{avg-k}$ gating and learned gating converge, as indicated by the lower halves of the plots. (2) $\operatorname{avg-k}$ gating is more stable than learned gating (Figure 6 a vs. b). This might mitigate data inefficiency from gating inconsistencies across different stages of training (Zuo et al., 2022). (3) EMoE, with its heterogeneous experts, exhibits better load balancing than GMoE (Figure 6 b vs. c). In GMoE, all experts share identical initialization, whereas in EMoE, the experts are derived from FFNs with implicit modularity. This also suggests a good initialization can facilitate MoEs learning similar to EvoMoE (Nie et al., 2021).
@@ 5.2 Ablation studies
Performance Across Different Training set Volumes. Previous research has indicated that modular architectures offer improved data efficiency (Bengio et al., 2020). Therefore, we conducted experiments with GPT2-XL on six tasks using varying proportions of original training data, and the results for all tasks are presented in Figure 7. It can be observed that EMoE consistently outperforms the dense across different data factions. EMoE achieves superior results even when using less than 20% of the data. On SST2, only using 50% data, EMoE shows comparable performance to the dense. More details can be found in the Appendix A.4.2 Table 15. This further underscores the benefits of incorporating modular structures.
6 Conclusion, Limitations and Future Works
Conclusion: In this work, we validate that exploiting the emergent modular structures in dense pre-trained transformers improves downstream task ID and OOD performances. One possible reason is the modular structure can alleviate negative transfer effects presented in the pre-trained model. We hope our findings could deepen the understanding of neural networks’ modularity, further helping the community develop more sophisticated modular neural architectures and utilizing existing pre-trained models. Limitations: We would like to point out that EMoE is not designed to outperform the SOTA tuning methods and OOD generalization algorithms. Instead, the goal is to demonstrate that dense pre-trained transformers can benefit from emergent modular structures. Our research findings have not been validated on more challenging tasks (e.g., Mathematical Reasoning (Imani et al., 2023)) or foundation models like LLaMA (Touvron et al., 2023) due to computational constraints. While our evaluation was primarily conducted on models with a maximum parameter count of 1.5B, it is evident that across the spectrum of parameter sizes, ranging from 22M to 1.5B, EMoE consistently yields improvements, showing the potential to scale up. Future works: Subsequent research could delve into better methods for modular decomposition and investigate the modularity in foundation models. Benefiting from EMoE’s ability to avoid negative transfer, future researchers can also exploit the emergent modularity in settings such as multi-task learning and continual learning. Moreover, it is observed that modularity structures emergent and remain stable after approximately one-fourth of the pre-training phase (Li et al., 2022; Zhang et al., 2023). Future investigations could investigate leveraging the modularity of dense models during the pre-training process.
7 Reproducibility
In the Experiment Configuration Section 4, we first introduce the additional configurations introduced by EMoE relative to the Dense Model, as well as the baseline settings. Subsequently, we provide high-level experiment settings for each experimental group (Domainbed at 4.1, full fine-tuning GLUE tasks at section 2, LoRA tuning at  3). In Appendix A.2, we provide detailed information on the datasets used, the codebase, and the evaluation metrics. We also outline more general configurations, such as learning rates and batch sizes for the respective tasks. Beyond the main paper, we include the original code for the experiments and log files for certain experimental results in the supplementary materials. Each section of code includes a README.md file that explains the experimental settings required to replicate the results.
Appendix A Appendix
A.1 Compare with Moefication
Table 5: Comparison between EMoE and Moefication
Aspect EMoE Moefication
Motivation
Modularity
Sparsity
Goal
Transfer learning effective
Inference efficiency
MoEs Layers
Typically 1 or 2 layers
Applies changes across all layers
Scenarios
Down-stream tasks
Original task
Results
Significant improvement without adding parameters
Improved inference efficiency but performance drop
Though adopting the MoEs construction method from MoEfication (Zhang et al., 2022b), EMoE has a distinct motivation, leading to different model configurations and testing scenarios. MoEfication applies the above method to each layer of fine-tuned models, hoping to boost inference efficiency by discarding inactivated neurons. In sharp contrast, we treat the decomposed experts as heterogeneous modules and explore the benefits of such modularity in downstream tasks. Accordingly, We only introduce EMoEs in certain layers according to GMoE, since efficiency is not our primary concern. These differences are summarized in Table 5.
A.2 Datasets and Evaluation Metrics
A.2.1 Domainbed
@@
The four datasets (PACS (Li et al., 2017), VLCS (Fang et al., 2013), Office-Home (Venkateswara et al., 2017), and Terra Incognita (Beery et al., 2018)) are selected from Domainbed. Each dataset comprises 4 distinct domains(PACS: {art, cartoons, photos, sketches}, VLCS: {Caltech101, LabelMe, SUN09, VOC2007}, Office-Home: {art,clipart, product, real}, and Terra Incognita: {L100, L38, L48, L46}). Within a single training, one or two domains’ data are sequentially designated for OOD evaluation. For example, when training on PACS, {art, cartoons} could be selected as ID training data, while {photos, sketches} are designated for OOD testing. This configuration results in a total of $C_{4}^{2}+C_{4}^{1}=10$ distinct training processes within each dataset. Suppose there are $d_{tr}$ ID domains, “Train-validation” means selecting OOD test checkpoints based on ID accuracies from the validation subsets of all $d_{tr}$ ID domains; “Leave-one-domain-out” means leaving one selected ID domain as a validation set, doing training on $d_{tr}-1$ domains; “Test-domain” means selection based on limited accesses to test domains and selecting based on these results. The final results are aggregated with selection criteria provided by Domainbed [^1] https://github.com/facebookresearch/DomainBed. As a result, even a variation of 0.1 in the benchmark outcomes signifies a significant improvement.
@@
A.2.2 GLUE
Each task involves one to four OOD tasks from GLUE-X (Yang et al., 2023), resulting in 13 OOD tasks in total. To illustrate, consider the Sentiment Analysis task: we first fine-tune models on SST-2 from GLUE and report the validation results as ID performance, then use the test data from IMDB (Maas et al., 2011), Yelp (Zhang et al., 2015), Amazon (Kaushik et al., 2020) and Flipkart (Vaghani & Thummar, 2023) from GLUE-X for OOD testing.
In the full fine-tuning, to ensure convergence and reduce randomness, we train all models 10 epochs across 3 random seeds on each task. Each experiment does a hyper-parameter search on learning rates on [2e-5, 3e-5, 5e-5] as suggested by BERT (Devlin et al., 2019). The training batch size is 32. In the LoRA tuning, following LoRA (Hu et al., 2022) that tunes models with more epochs and larger learning rates than standard full fine-tuning, all models are trained 20 epochs on small and medium datasets and 5 epochs on large ones (like QNLI, MNLI, QQP). The learning rate is searched in [2e-4, 3e-4, 5e-4]. All methods are implemented with LoRA_rank=8 and LoRA_alpha=16. The training batch size is 16 due to a larger model size. Other settings like max_lengt following the codebase from hugging face [^2] https://github.com/huggingface/transformers/tree/main/examples/pytorch/text-classification. After training on GLUE, we directly test the selected models on GLUE-X with the data from the official repo [^3] https://github.com/YangLinyi/GLUE-X.

### v2

@@ B.1 Full Fine-tuning Performance
We evaluate EMoE’s ID performance on 5 GLUE tasks. We also include an additional baseline noisy tuning (Wu et al., 2022), which improves adaptation for free by adding uniform distribution noise proportional to the standard deviation of the pre-trained weights before tuning. According to Table 6: (1) On average, EMoE and EMoE-learn outperform other baselines. (2) Among the two methods that do not introduce additional parameters, EMoE significantly outperforms noisy tuning. (3) EMoE provides stable improvements over baselines across different settings, demonstrating its generality.
Table 7: Instruction full-tuned Llama2-7B’s MMLU scores. Times are wall-clock computation times. The term ’times’ in the subsequent tables refers to the same concept. ’w/o’ refers to ’without’.
Experts (N) topk (k) MMLU score times (s) FLOPS ($10^{16}$)
w/o tuning - 46.79 - -
full tuning - 46.5 4988 8.97
64 16 48.08 5036 9.12
64 32 47.44 5041 9.24
B.2 Mult-tasks Setting
Table 8: T5-base ID performances. All tasks are from SuperGLUE so we omit the prefix "SuperGLUE" for each tasks.
Experts (N) topk (k) boolq cb wic wsc.fixed rte copa test avg
Baseline 82.14 85.71 65.83 34.62 74.10 52.00 65.73
8 2 80.67 89.29 65.52 36.54 79.86 56.00 67.98
16 4 81.16 89.29 68.34 51.92 74.10 44.00 68.14
32 8 80.12 78.57 70.85 63.46 82.73 64.00 73.29
32 16 81.04 75.00 72.41 57.69 78.42 54.00 69.76
Table 9: T5-base OOD performances. ’SG’ refers to ’SuperGLUE’.
Experts (N) topk (k) mnli qnli wnli rte SG-rte SG-cb OOD avg
Baseline 86.2 92.42 50.00 61.87 62.59 32.14 51.65
8 2 86.27 92.18 50.00 64.03 62.59 32.14 52.19
16 2 86.44 92.59 52.78 64.03 56.83 39.29 53.23
32 8 86.56 92.49 58.33 64.03 61.87 28.57 53.20
In our analysis 4, we have identified that the improvement brought by EMoE is likely associated with mitigating negative transfer. Inspired by this, we choose a multi-task learning setting where negative transfer might be more pronounced. We adopt the codebase [^3] https://github.com/AkariAsai/ATTEMPT from ATTEMPT (Asai et al., 2022). For the in-domain (ID) scenario, we follow the settings outlined in ATTEMPT and select six tasks from the Super-GLUE benchmark (Wang et al., 2019a). For the out-of-domain (OOD) scenario, we take two larger natural language inference (NLI) datasets MNLI and QNLI, from GLUE as our ID training data. We subsequently conducted direct testing on four additional NLI datasets from different domains. All hyperparameters unrelated to MoEs are kept consistent with the baseline, and we have listed the MoEs-related hyperparameters in the result table. Our observations are as follows: 1. EMoE exhibits a substantial improvement compared to the baseline. In the in-domain (ID) setting, the highest improvement reached 7.56, even considering the average performance across the six tasks. In the OOD setting, the highest average OOD result across the four datasets improved by 1.58. 2. Across various settings of N and K, EMoE consistently outperforms the vanilla fone-tuning. Within the hyperparameter search space specified in our paper, EMoE consistently improves at least 2 points over the baseline in the ID setting. This emphasizes the effectiveness of EMoE and EMoE’s robustness to the explored hyperparameter range.
B.3 Performance Across Different Training set Volumes
@@
Previous research has indicated that modular architectures offer improved data efficiency (Bengio et al., 2020). Therefore, we conducted experiments with GPT2-XL on six tasks using varying proportions of original training data, and the results for all tasks are presented in Figure 7. It can be observed that EMoE consistently outperforms the standard across different data factions. EMoE achieves superior results even when using less than 20% of the data. On SST2, only using 50% data, EMoE shows comparable performance to the standard. More details can be found in the Appendix F.2 Table 20. This further underscores the benefits of incorporating modular structures.
Appendix C Comparison between EMoE, MoEfication, and GMoE
Table 10: Comparison between EMoE, MoEfication, and GMoE
Aspect EMoE MoEfication GMoE
Research Problem Exploit the emergent modularity during fine-tuning pre-trained transformers. Approximate FFNs with sparse MoE to improve inference efficiency. Validate the OOD improvements brought by Sparse MoE architectures.
Methods Split FFNs Split FFNs Copy of FFNs
Practicality No additional trainable parameters. Experts can be recomposed into the standard model so that models can be deployed as a standard model. May need re-training on the original task. May suffer from inference latency owing to the specific implementation of MoE architectures. Additional trainable parameters are introduced. May suffer from inference latency owing to the specific implementation of MoE architecture.
Contribution Significant general improvement without adding parameters and not depending on the specific implementation. Improved inference efficiency (depending on the specific implementation of MoE), but performance drop. Significant OOD improvement with additional parameters and specific implementation.
Leveraging modular designing into neural networks has various advantages (Pfeiffer et al., 2023; Chowdhery et al., 2022; Chen et al., 2023a), including superior generalization abilities (Goyal et al., 2021; Li et al., 2023). Most MNNs are explicitly modular. Among them, MoEs (Szymanski and Lemmon, 1993) is currently a standard architecture for developing MNNs (Shen et al., 2023b; Shazeer et al., 2017; Fedus et al., 2022; Zhang et al., 2022a). Apart from explicit MNNs,  Hod et al. (2021); Csordás et al. (2021) explore emergent modular structures in CNNs and LSTMs. Some recent works (Zhang et al., 2022b; Li et al., 2022) focus on the sparsity of more complicated pre-trained transformers. Based on their observations,  Zhang et al. (2023) recently explore modularity in pre-trained transformer FFNs by employing handpicked semantic and knowledge-intensive tasks to probe the modular nature of pre-trained transformers.
Having observed that FFN layers in pre-trained Transformers are sparsely activated (many neurons are unused for inputs), MoEfication splits transforms FFNs into a sparse MoE, aiming to approximate the functionality of the original FFNs to reduce the computational cost, further improving inference efficiency. Besides, the GMoE makes multiple replicates of the original FFN layer and introduces a learned gate to form a MoE architecture. They claim that such architecture could improve OOD performance from their theoretical perspective of algorithmic alignment framework. MoEfication and GMoE do not touch on how emergent modularity influences the training stage. The table below illustrates a detailed comparison of these works. These differences are summarized in Table 10.
Appendix D Datasets and Evaluation Metrics
D.0.1 Domainbed
@@
The four datasets (PACS (Li et al., 2017), VLCS (Fang et al., 2013), Office-Home (Venkateswara et al., 2017), and Terra Incognita (Beery et al., 2018)) are selected from Domainbed. Each dataset comprises 4 distinct domains(PACS: {art, cartoons, photos, sketches}, VLCS: {Caltech101, LabelMe, SUN09, VOC2007}, Office-Home: {art,clipart, product, real}, and Terra Incognita: {L100, L38, L48, L46}). One or two domains’ data are sequentially designated within a single training for OOD evaluation. For example, when training on PACS, {art, cartoons} could be selected as ID training data, while {photos, sketches} are designated for OOD testing. This configuration results in $C_{4}^{2}+C_{4}^{1}=10$ distinct training processes within each dataset. Suppose there are $d_{tr}$ ID domains, “Train-validation” means selecting OOD test checkpoints based on ID accuracies from the validation subsets of all $d_{tr}$ ID domains; “Leave-one-domain-out” means leaving one selected ID domain as a validation set, doing training on $d_{tr}-1$ domains; “Test-domain” means selection based on limited access to test domains and selection based on these results. The final results are aggregated with selection criteria provided by Domainbed [^4] https://github.com/facebookresearch/DomainBed. As a result, even a variation of 0.1 in the benchmark outcomes signifies a significant improvement.
@@
D.1 GLUE
Each task involves one to four OOD tasks from GLUE-X (Yang et al., 2023), resulting in 13 OOD tasks in total. To illustrate, consider the Sentiment Analysis task: we first fine-tune models on SST-2 from GLUE and report the validation results as ID performance, then use the test data from IMDB (Maas et al., 2011), Yelp (Zhang et al., 2015), Amazon (Kaushik et al., 2020) and Flipkart (Vaghani and Thummar, 2023) from GLUE-X for OOD testing.
In the full fine-tuning, to ensure convergence and reduce randomness, we train all models 10 epochs across 3 random seeds on each task. Each experiment does a hyper-parameter search on learning rates on [2e-5, 3e-5, 5e-5] as suggested by BERT (Devlin et al., 2019). The training batch size is 32. In the LoRA tuning, following LoRA (Hu et al., 2022) that tunes models with more epochs and larger learning rates than standard full fine-tuning, all models are trained 20 epochs on small and medium datasets and 5 epochs on large ones (like QNLI, MNLI, QQP). The learning rate is searched in [2e-4, 3e-4, 5e-4]. All methods are implemented with LoRA_rank=8 and LoRA_alpha=16. The training batch size is 16 due to a larger model size. Other settings like max_lengt following the codebase from hugging face [^5] https://github.com/huggingface/transformers. After training on GLUE, we directly test the selected models on GLUE-X with the data from the official repo [^6] https://github.com/YangLinyi/GLUE-X.